OpenStudy (anonymous):

10x - 10 -7x > with line under it 3x - 2 x -8 x 8 all real numbers no solution

OpenStudy (anonymous):

You mean \[10x - 10 -7x \geq 3x-2\] and you want to find a restriction on \[x.\] Ok well tidy up the inequality: \[3x-10 \geq 3x-2\] since there is a 3x on both sides, these go away to give simply: \[-10 \geq -2\] but this is VERY BAD because -10 < -2. So the inequality is inconsistent, and so there cannot exist an x which satisfies it. Therefore there is no solution.